OpenStudy (anonymous):

Two planes have left the L.A. airport. One plane is 30 miles west of the airport and has an altitude of 4 miles, while the other plane is 40 miles east of the airport and has an altitude of 2 miles. What is the distance between the planes?

OpenStudy (anonymous):

Did you draw this out?

OpenStudy (anonymous):

no cuz idk how

OpenStudy (anonymous):

I know the answer is a square root

OpenStudy (anonymous):

This help

OpenStudy (anonymous):

not really

OpenStudy (anonymous):



OpenStudy (anonymous):

70?

OpenStudy (anonymous):

wait no

OpenStudy (anonymous):

160

OpenStudy (anonymous):

no

OpenStudy (anonymous):

 You need to determine the lengths of the two hypotenuses I have put ? marks on

OpenStudy (anonymous):

how do I figure it out

OpenStudy (anonymous):

30 times 4=120?

OpenStudy (anonymous):

They are right triangles? Do you know the Pythagorean formula for right triangles?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

a^2+b^2=c^2

OpenStudy (anonymous):

4^2+30^2=916 c=sq rt 916=30.2655 2^2+40^2=1604 sq rt of 1604=40.05